RO-VIBRATIONAL ANALYSIS OF THE ν_{4}, ν_{6} and ν_{3} BANDS OF THIOFORMALDEHYDE: EXAMPLE OF A MASSIVE Z-TYPE CORIOLIS RESONANCE

\underline{W. J. LAFFERTY}, Optical Technology Division, NIST Gaithersburg, MD 20899, USA; J.-M. FLAUD, A. PERRIN, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Universités Paris 12 et 7, 61 Av du Général de Gaulle, 94010 Créteil Cedex France; H. BECKERS, Y.S. KIM, H. WILLNER, Anorg. Chemistry, University of Wuppertal, D-42119 Wuppertal, GERMANY.

The infrared spectrum of thioformaldehyde (CH_{2}S) is of interest since this species is one of the myriad of molecules found in the interstellar space. From a spectroscopic point of view, it is of even more interest since the two lowest vibrational modes, the in-plane rocking mode, ν_{6} (B_{2}), and the out-of-plane wagging mode, ν_{4} (B_{1}), fall at 990.18 and 991.02 cm^{−1} respectively. This separation of only 0.84 cm^{−1} leads to a massive z-type Coriolis resonance where many of the rotational levels of each of the two vibrational states are mixed nearly 50% with each other. To make the situation even more interesting the C=S stretching vibration, ν_{3}, with A1 symmetry occurs nearby at 1059 cm^{−1}. This vibrational level also interacts with the two low frequency modes which complicates the assignment and analysis. CH_{2}S was produced by low pressure thermolysis of a gas flow of C_{3}H_{5}SCH_{3}/Ar (560^{∘} C) and CH_{3}SCl/Ar (1150^{∘} C) in the entrance of the multipath white cell (optical path length 32 m). At a total pressure of 0.15 mbar, 40 scans were recorded for the range 750 to 1400 cm^{−1} on a Bruker HR120 TFIR spectrometer at a resolution of 0.005 cm^{−1} (maximum optical path difference). The initial line assignment was not straightforward. There are strong series apparent in the spectrum, but the features expected for a b-type and c-type bands were not obvious near the band center. The centers of these three vibrations have been determined from medium resolution FT spectra^{a} as well as laser Stark measurements^{b}. An initial calculation was made using this information as well as guessed values for the band intensities. This permitted the identification of several low K_{a} series. Finally after numerous iterations, the transitions in the spectrum were identified leading to an excellent set of ro-vibrational constants.

aC. Clouthier, D.C. Moule, D.A. Ramsay and F.W. Birss, Can. J. Phys 60, 1212 (1982)

bD.J. Bedwell and G. Duxbury, J. Mol. Spectrosc. 64, 531 (1980)